Circular No. 7169

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999cd IN NGC 3646 M. Papenkova and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report their discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova in NGC 3646. The new object was discovered on unfiltered images taken on May 14.224 and 14.279 UT (mag ∼ 17.9) and was confirmed on images taken on May 15.175 (mag ∼ 17.6). SN 1999cd is located at α = 11^{h}21^{m}42^{s}.89, δ = +20^{o}09′24′′.1 (equinox 2000.0), which is ∼ 3′′.7 west and 46′′.4 south of the galaxy nucleus. A KAIT image of the same field taken on May 9.2 shows nothing at the position of SN 1999cd (limiting mag 19.0). SN 1989N, a type-II supernova, also occurred in NGC 3646 (peak mag ∼ 14).

SUPERNOVA 1999cc IN NGC 6038 P. Garnavich, S. Jha, R. Kirshner, and P. Challis, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1999cc (cf. IAUC 7163) was obtained by P. Berlind with the 1.5-m Tillinghast telescope on May 14.3 UT that shows this to be a type-Ia supernova before maximum brightness. A strong Si II absorption is observed at 628.9 nm, which implies a photospheric velocity of 12 300 km/s after correcting for the redshift of NGC 6038 (9500 km/s, derived from narrow emission in the spectrum).

SUPERNOVA 1999cb IN MARKARIAN 881 Garnavich et al. also write: “A spectrum of SN 1999cb (cf. IAUC

7163) was obtained by P. Berlind with the 1.5-m Tillinghast telescope on May 14.4 UT. The spectrum shows a deep, broad Na I 586-nm absorption and iron features in the blue, which suggest that this a type-Ia supernova about a month past maximum brightness.”

SUPERNOVA 1999by IN NGC 2841 CCD magnitudes by D. Hanžl, Brno, Czech Republic (cf. IAUC 7157,

7158): May 2.872 UT, V − R = +0.12 ± 0.03; 5.821, V = 13.71 ± 0.01, B − V = +0.32 ± 0.01, V − I = +0.28 ± 0.03, V − R = +0.19 ± 0.01; 7.049, V = 13.52 ± 0.02, B − V = +0.31 ± 0.1, V − I = +0.25 ± 0.02; 14.046, V = 13.16 ± 0.09, V − R = +0.3 ± 0.1.

1999 May 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green